# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v3

@@ 7.2 Sensitivity analysis for the signature-based detector
We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the lead-time versus alert-burden trade-off. Table 2 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median lead time over fatigue creatives.
@@
Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median lead (days)
@@ 7.2.1 Financial Impact Quantification
The true value of earlier detection becomes apparent when translated to operational units. The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel). Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and an absolute CTR shortfall of $\Delta CTR$ relative to a benchmark period, the implied opportunity cost per day is
$\text{Opportunity Cost per day}=I\,\Delta CTR\,CPC_{bench}.$
@@ 8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series, as discussed in Section 4. This makes it highly scalable and efficient. A typical analysis for a single creative with a year’s worth of daily data can be completed in a fraction of a second on standard hardware.
The methodology can be deployed in either a batch or a real-time setting. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an entire portfolio of active creatives, generating a summary report of assets that may be exhibiting fatigue. In a more advanced, real-time implementation, the algorithm could be integrated with a streaming data pipeline, allowing for the continuous monitoring of creative performance and the generation of automated alerts when a potential change point is detected. This would enable the fastest possible response to creative fatigue.
@@ 8.2 Managerial Implications
The primary implication of this work for marketing managers is the ability to move from a reactive, heuristic-based approach to creative management to a proactive, data-driven one. By providing a reliable and early warning system for creative fatigue, our methodology enables several key managerial actions.
First, it allows for more efficient budget allocation. Instead of continuing to spend on a creative that is no longer effective, managers can reallocate that budget to higher-performing assets or to the development of new creative. The wastage calculation provides a direct financial justification for these decisions.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives need to be replaced or updated most urgently, based on the magnitude of their performance decline and the associated financial wastage.
Third, it creates a more rigorous feedback loop for the creative development process itself. By analysing which types of creatives fatigue more quickly, marketing teams can gain insights into what resonates most with their audience over the long term, informing the development of more durable and effective advertising campaigns in the future.
@@ 9.1 Ethical Considerations
The methodology presented in this paper operates on aggregated, anonymised performance data. It does not require access to any personally identifiable information (PII) or individual user-level data, and as such, it aligns with modern data privacy principles such as those outlined in GDPR and other regulations. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
@@ 10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from a synthetic panel evaluation indicate that this signature-based approach can provide early warnings under an explicit operational onset definition, while highlighting a practical trade-off between early warning and alert burden. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of opportunity cost provides a bridge from statistical outputs to operational prioritisation.
The practical implications of this methodology for marketing practitioners are significant. The automated and scalable nature of the algorithm allows for the continuous monitoring of a large portfolio of creative assets, something that is infeasible to do manually. This enables marketing and media teams to move from a reactive to a proactive stance on creative management. Rather than waiting for performance to decline substantially, they can be alerted to the onset of fatigue in near real-time. The wastage metric provides a clear, data-driven rationale for when to retire an underperforming creative and reallocate its budget, thereby improving the overall efficiency and return on investment of advertising campaigns.
@@ Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path.
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}$
@@
These second-order terms capture the area swept out by the path in the $ij$-plane. Higher-order terms capture more complex geometric properties.
A key result from rough path theory, Chen’s theorem, states that the signature uniquely determines the path up to tree-like deformations. This means that if two paths have the same signature, they are essentially equivalent from a geometric perspective. This uniqueness property is what makes the signature such a powerful feature set for describing path-like data. In practice, we use a truncated signature, which provides a finite-dimensional approximation that is sufficient for most machine learning applications. Our library provides an efficient and numerically stable implementation for computing these truncated signatures.
@@ Appendix B Algorithm Parameters
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the nature of the time series being analysed. For daily data from typical digital advertising campaigns, we have found that a window size between 7 and 14 days provides a good balance between sensitivity and stability.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives. It determines how large the distance between two consecutive signatures must be to be considered a statistically significant change. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. A value of $k$ between 1.5 and 2.5 has been found to work well in practice, providing a reasonable balance for typical performance marketing data. The optimal choice can be fine-tuned using a validation set if historical data with known change points is available.
Appendix C Mathematical Foundations and Theoretical Analysis
This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory. For readability, we present proof sketches rather than fully detailed proofs. We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined. To avoid notational ambiguity, we use $m$ to denote the path dimension and $d$ to denote the signature truncation depth.
@@ C.1 Signature distance sensitivity
Theorem 1 (Signature sensitivity to path changes (proof sketch)).
Let $X,Y:[0,T]\to\mathbb{R}^{2}$ be two continuous paths representing performance trajectories, with signatures $S(X)$ and $S(Y)$ truncated at depth $d$. If there exists a time interval $t_{1},t_{2}]\subset[0,T]$ where the paths exhibit different geometric properties (trend, volatility, or curvature), then:
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|,$
where $C>0$ depends on regularity assumptions and the truncation depth $d$.
Proof.
The signature captures iterated integrals of the path increments. For the first level:
@@
This directly captures the difference in total displacement. For the second level, which captures signed area:
@@
This term is sensitive to differences in the path’s oscillatory behaviour and volatility. By the Chen-Strichartz formula, the signature norm satisfies:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|\geq\|X(T)-X(0)-(Y(T)-Y(0))\|$
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. A rigorous statement requires explicit norms and assumptions (e.g., bounded $p$-variation) and follows from continuity estimates for truncated signatures. ∎
C.2 Theoretical Analysis of Detection Power
Theorem 2 (Detection power (heuristic bound)).
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Let the process have different distributions before and after $\tau$:
@@ C.2 Theoretical Analysis of Detection Power
For window size $w$ and threshold $\theta=\mu_{D}+k\sigma_{D}$, the detection power (probability of correctly identifying the change) is:
$\text{Power}=P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{w\cdot\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}{2\sigma^{2}}\right)$
where $\text{KL}(\cdot||\cdot)$ is the Kullback-Leibler divergence and $\sigma^{2}$ is the variance of the signature distance under the null hypothesis.
Proof.
Under $H_{1}$ (change present), windows $W_{1}$ and $W_{2}$ straddling the change point contain samples from different distributions. The signature distance $D$ between these windows can be bounded using the data processing inequality:
$D^{2}=\|S(X_{W_{1}})-S(X_{W_{2}})\|^{2}\geq w\cdot d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})^{2}$
where $d_{\text{TV}}$ is the total variation distance.
By Pinsker’s inequality:
$d_{\text{TV}}(\mathcal{F}_{1},\mathcal{F}_{2})\geq\sqrt{\frac{1}{2}\text{KL}(\mathcal{F}_{1}||\mathcal{F}_{2})}$
Under $H_{0}$, the signature distances concentrate around their mean; a complete derivation depends on dependence assumptions in the time series and on the distribution of the signature estimator. Using concentration-style reasoning:
$P(D>\theta|H_{1})\geq 1-\exp\left(-\frac{(\mathbb{E}[D|H_{1}]-\theta)^{2}}{2\sigma^{2}}\right)$
Substituting lower bounds on $\mathbb{E}[D|H_{1}]$ yields the stated qualitative dependence. The intent is to highlight that detection improves with larger window sizes and with greater separation between pre- and post-change regimes. ∎
C.3 Convergence Properties as Window Size Increases
Theorem 3 (Asymptotic consistency (informal)).
Let $w$ be the window size and $n$ be the total number of observations. As $w\to\infty$ with $w=o(n)$, the signature-based change point estimator $\hat{\tau}_{w}$ converges to the true change point $\tau$ in probability:
$P(|\hat{\tau}_{w}-\tau|>\epsilon)\to 0\quad\text{as }w\to\infty$
for any $\epsilon>0$.
Proof.
@@ C.3 Convergence Properties as Window Size Increases
Under regularity conditions (e.g., bounded $p$-variation and appropriate mixing), this process satisfies:
  1. 1.
    For $t$ far from $\tau$: $\mathbb{E}[D_{t}(w)]=O(w^{-1/2})$ (by CLT for signatures)
  2. 2.
    For $|t-\tau|<w$: $\mathbb{E}[D_{t}(w)]\geq c\cdot\min(|t-\tau|/w,1)$ for some $c>0$
The change point estimator is:
$\hat{\tau}_{w}=\arg\max_{w<t<n-w}D_{t}(w)$
By the continuous mapping theorem and the fact that $D_{t}(w)$ has a unique maximum near $\tau$ (with high probability), we have:
$\hat{\tau}_{w}\to\tau\quad\text{in probability}$
The rate of convergence depends on additional assumptions and is stated here for intuition only. Larger window sizes provide more stable estimation at the cost of reduced temporal resolution. ∎
C.4 Computational Complexity Analysis
Lemma 1 (Signature computation complexity (order of growth)).

### v4

@@ 7.4 Sensitivity analysis for the signature-based detector
We next vary the signature detector’s window size $w_{d}$ and threshold multiplier $k$ to characterise the alert-timing versus alert-burden trade-off. Table 17 reports the number of fatigue creatives alerted (out of 6), burden (out of 11), and the median signed detection offset relative to the fatigue reference date over alerted fatigue creatives only. The offset is $L=\tau-\hat{\tau}$, so positive values indicate a marker before the reference date and negative values indicate a marker after it.
@@
Window $w_{d}$ Threshold $k$ Alerts on fatigue Burden / Median signed $L$
@@ 7.4.1 Financial Impact Quantification
The value of earlier review can be translated into operational units. The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel). Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and a nonnegative CTR shortfall $\Delta CTR\geq 0$ relative to a benchmark period, the implied spend-equivalent proxy per day is
$\text{Spend-equivalent proxy per day}=I\,\Delta CTR\,CPC_{bench}.$
@@ 8.1 Implementation Considerations
The signature-based detection algorithm is designed to be integrated into existing marketing analytics workflows. In terms of computational requirements, the algorithm exhibits linear complexity with respect to the length of the time series for fixed path dimension and signature depth, as discussed in Section 4. The runtime smoke benchmark in Table 16 supports the feasibility of batch portfolio screening on synthetic trajectories, although it does not replace an end-to-end system benchmark.
The methodology is most directly supported as a batch or periodic screening tool. In a batch processing mode, the analysis can be run periodically (e.g., daily or weekly) across an active creative portfolio, generating a summary report of assets that may be exhibiting fatigue. A deployable online protocol should be specified before operational use. A conservative version is: require at least 45 observed calibration days, compute signatures on the Beta–Binomial smoothed CTR series, estimate the distance threshold only from past distances, require a directional post-window mean drop for fatigue-specific alerts, suppress repeat alerts for a fixed cooldown period, and tune the threshold using an explicit review-cost or alert-budget objective such as Table 10. The pure shape-only signature detector can still be useful as a generic trajectory-change screen, but the injected benchmark indicates that fatigue monitoring should use the level-aware rule. Real-time implementation would still require exposure-aware noise handling, alert-budget controls, and portfolio-level runtime validation before deployment claims could be made.
@@ 8.2 Managerial Implications
The primary implication of this work for marketing managers is a more structured screening input for periodic creative review, rather than a fully validated replacement for existing heuristic decision rules. By providing a screening signal for possible creative fatigue, our methodology enables several managerial actions.
First, it can support more efficient budget allocation discussions. Instead of relying only on manual review or monthly averages, managers can inspect creatives whose trajectories have changed and decide whether they warrant further diagnosis before any budget reallocation or creative refresh. The wastage proxy provides a prioritisation signal, not a direct causal financial justification.
Second, it provides a quantitative basis for the creative refresh cycle. The output of the analysis can be used to prioritise which creatives should be reviewed most urgently, based on the magnitude of their performance decline and the associated benchmark-relative opportunity proxy.
Third, it can create a more rigorous feedback loop for the creative development process itself. By tracking which types of creatives are repeatedly flagged for review, marketing teams can generate hypotheses about durability and audience response that should be checked against campaign context before being treated as substantive audience-preference findings.
@@ 9.1 Ethical Considerations
The methodology presented in this paper can operate on aggregated performance data and does not require individual user-level data. This is privacy-favourable compared with user-level tracking, but it does not by itself establish GDPR compliance; compliance depends on the broader data collection, consent, retention, and governance context in which the method is deployed. The analysis focuses on the performance of the creative asset itself, not on the behaviour of individual users. However, any application of performance measurement could indirectly lead to algorithmic bias. For example, if the system consistently flags creatives targeted at a specific demographic as fatiguing more quickly, it could lead to a reduction in advertising shown to that group. It is therefore important that practitioners use this tool as one input into a broader decision-making process, and remain mindful of the potential for such second-order effects.
@@ 10 Conclusion
In this paper, we have introduced a methodology for screening advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from a synthetic panel evaluation and stochastic injected-changepoint benchmark indicate that this signature-based approach can provide a structured trajectory-change signal, while highlighting a practical trade-off between event recovery and alert burden. In the injected benchmark, the original shape-only signature detector is not sufficient for fatigue dominated by absolute level deterioration; within that benchmark, the most defensible candidate workflow among the evaluated online signature variants is the level-aware signature screen with explicit directional gating and threshold tuning. The subsequent benchmark-relative opportunity proxy provides a bridge from statistical outputs to operational prioritisation.
The practical implication of this methodology is a structured screening workflow for creative review. The detector can highlight candidate creatives whose performance trajectory has changed, and the benchmark-relative wastage proxy can help prioritise analyst attention. The present synthetic evidence does not show that the method should autonomously retire creatives or reallocate budgets; rather, it supports use as one input into a broader decision process. Future work within the same data-governance constraint should extend the baseline set further (for example PELT, Bayesian online changepoint detection, and exposure-weighted changepoint rules), evaluate broader synthetic nuisance scenarios, report full alert-budget curves, and improve synthetic calibration diagnostics.
@@ Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of an appropriately regular path. For a bounded-variation path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$. More irregular continuous paths require an appropriate rough-path enhancement before these iterated integrals are well defined.
The first level of the signature, $S^{1}(X)$, is a vector representing the total displacement of the path. For bounded-variation paths,
$S^{1}(X)_{i}=\int_{0}^{1}dX_{t}^{i}=X_{1}^{i}-X_{0}^{i}.$
@@
These second-order terms capture ordering information in the path increments. The antisymmetric component, $\frac{1}{2}(S^{2}_{ij}-S^{2}_{ji})$, corresponds to signed oriented area in the $(i,j)$ plane; the full second level should not be interpreted simply as planar area. Higher-order terms capture more complex geometric properties.
Chen’s identity states that signatures are multiplicative under path concatenation, which is central to their algebraic usefulness. Separately, uniqueness results for bounded-variation paths with a fixed starting point show that the full signature determines the path up to tree-like equivalence under suitable conditions (Hambly and Lyons, 2010). In practice, we use a truncated signature, which provides a finite-dimensional approximation to this infinite object. Truncation generally loses injectivity: many distinct paths can share the same low-order signature terms. The empirical question is therefore whether the chosen truncation depth preserves the path features relevant to advertising-performance monitoring.
@@ Appendix B Algorithm Parameters
The window size $w$ determines the granularity of the analysis. A smaller window size will be more sensitive to short-term fluctuations and can detect changes more quickly, but it may also be more susceptible to noise and false positives. A larger window size will provide a more stable estimate of the path signature, making it more robust to noise, but it will be slower to detect changes and may miss short-lived events. The optimal choice of $w$ depends on the cadence, volatility, and review objective of the time series being analysed. The experiments in this paper use daily windows in the 7–14 day range as exploratory operating points, not as universal defaults.
The sensitivity threshold $k$ controls the trade-off between false positives and false negatives through $D_{i}>\mu_{D}+k\sigma_{D}$, where $\mu_{D}$ and $\sigma_{D}$ are computed from the chosen reference distribution of adjacent-window distances. In the retrospective panel experiment this distribution is the full creative-specific distance series; in an online detector it must be a past-only calibration distribution. A smaller value of $k$ will make the detector more sensitive, leading to more change points being identified, but potentially including more false positives. A larger value of $k$ will make the detector more conservative, reducing the number of false positives but increasing the risk of missing true changes. Values between 1.5 and 2.5 are explored in this paper as operating points, not as universal defaults. The optimal choice can be fine-tuned using a validation set, alert budget, or historical data with known change points if available.
Appendix C Mathematical Foundations and Theoretical Remarks
This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory. The statements below are not used as formal guarantees for the empirical results. They are intended to clarify the intuition behind the representation and to state the assumptions that would be needed for stronger theory. We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined. To avoid notational ambiguity, we use $m$ to denote the path dimension and $d$ to denote the signature truncation depth.
@@ C.1 Signature distance sensitivity
The signature captures iterated integrals of path increments. For the first level:
@@
This directly captures differences in total displacement. For the second level:
@@
For multidimensional embeddings, second-level cross terms encode ordering information through interactions between channels. In one-dimensional paths, the second level reduces to a function of endpoint displacement,
$S^{2}(X)^{11}_{[t_{1},t_{2}]}=\frac{1}{2}\{X(t_{2})-X(t_{1})\}^{2},$
and therefore does not by itself capture oscillation. If the truncated signature is equipped with a Euclidean direct-sum norm whose level-1 block has unit weight, the distance dominates the first-level displacement difference:
$\|S(X)-S(Y)\|\geq\|S^{1}(X)-S^{1}(Y)\|=\|X(T)-X(0)-(Y(T)-Y(0))\|.$
Under other equivalent finite-dimensional norms, the same statement holds only up to a norm-dependent constant. Higher-order and cross-channel terms can distinguish some paths with similar endpoints but different shapes, but a general deterministic lower bound in terms of the path sup-norm requires additional assumptions and is not claimed here.
C.2 Depth-1 relation to adjacent-window statistics
It is useful to spell out what the level-aware construction reduces to at depth 1. Consider two adjacent windows represented as paths
$X^{L}(u)=(u,z_{L}(u)),\qquad X^{R}(u)=(u,z_{R}(u)),\qquad u\in[0,1],$
where $z$ is the calibration-standardised smoothed CTR. Their first-level signatures are endpoint increments:
$S^{1}(X^{L})=(1,z_{L}(1)-z_{L}(0)),\qquad S^{1}(X^{R})=(1,z_{R}(1)-z_{R}(0)).$
The time-coordinate increments cancel in the adjacent-window distance, so the depth-1 level-aware signature distance is
$\|S^{1}(X^{L})-S^{1}(X^{R})\|_{2}=|(z_{L}(1)-z_{L}(0))-(z_{R}(1)-z_{R}(0))|.$
Thus, for affine windows $z_{L}(u)=a_{L}+b_{L}u$ and $z_{R}(u)=a_{R}+b_{R}u$, the depth-1 signature distance reduces to $|b_{L}-b_{R}|$: it compares endpoint slopes over the two windows, not their mean level difference. The fatigue-specific rule in Algorithm 1 therefore combines a signature displacement or shape-change statistic with a separate directional mean-drop gate,
$G_{\tau}=\bar{z}_{L(\tau)}-\bar{z}_{R(\tau)}.$
This gate is close in spirit to adjacent-window mean-drop or MOSUM-style summaries, while the higher signature levels add terms that can respond to ordering and curvature within the windows. If one instead augments the path with a cumulative-level coordinate,
$Q(u)=\int_{0}^{u}z(v)\,dv,$
then the first-level displacement in $Q$ equals the window mean over $[0,1]$. In that alternative embedding, the depth-1 adjacent-window distance directly contains $|\bar{z}_{L}-\bar{z}_{R}|$. This shows that simple mean-drop detectors are recoverable as low-order signature features under an appropriate embedding, and it clarifies the incremental role of the higher-order path terms evaluated in the benchmark.
C.3 Detection power intuition
Consider a time series $\{Y_{t}\}$ with a change point at $\tau$. Suppose the process has different distributions before and after $\tau$:
@@ C.3 Detection power intuition
Under $H_{1}$ (change present), adjacent windows straddling the change point may have different expected path shapes, and therefore larger expected signature distances, than windows drawn from a stable regime. Under $H_{0}$, the distribution of signature distances depends on the sampling process, window overlap, exposure levels, and serial dependence. A complete power analysis would therefore require a specified stochastic model for impressions, clicks, and latent CTR dynamics. We do not claim a closed-form lower bound from KL divergence or total variation to a realised Euclidean signature distance; data-processing inequalities concern divergences between distributions after measurable transformations, not deterministic lower bounds on one sampled path.
Standard Pinsker’s inequality is
$d_{\mathrm{TV}}(P,Q)\leq\sqrt{\frac{1}{2}\mathrm{KL}(P\|Q)},$
or equivalently $\mathrm{KL}(P\|Q)\geq 2d_{\mathrm{TV}}(P,Q)^{2}$. This inequality can be useful in a fully specified distributional analysis, but it is not sufficient by itself to prove a lower bound for the detector used here.
C.4 Convergence Properties as Window Size Increases
Let $w$ be the window size and $n$ be the total number of observations. A plausible consistency target for a window-based estimator is localisation at the scale of the window, for example:
$P(|\hat{\tau}_{w}-\tau|/w>\epsilon)\to 0\quad\text{as }w\to\infty,\quad w=o(n).$
This scaled statement is weaker and more appropriate than exact recovery of an integer changepoint index.
@@ C.4 Convergence Properties as Window Size Increases
Under regularity conditions (e.g., bounded $p$-variation, appropriate mixing, and a unique separation of pre- and post-change regimes), one would need to show that this process is maximised near $\tau$ with high probability. A proof would also need to account for overlapping windows and threshold estimation. The present paper does not rely on such a theorem; it uses simulation evidence and reports alert burden explicitly.
C.5 Noise and exposure considerations
The signature is a continuous functional of the path under appropriate rough path metrics, so small perturbations of a path produce controlled perturbations of its signature. This continuity supports the use of signatures as stable path features. However, observed CTR has heteroskedastic binomial noise:
$\operatorname{Var}(C_{t}/I_{t}\mid I_{t},p_{t})=\frac{p_{t}(1-p_{t})}{I_{t}}.$
Consequently, min-max scaling of realised CTR paths does not remove the dependence of noise level on impression volume. Exposure-aware variants–for example signatures of $(t,\widehat{p}_{t},\log I_{t})$, distance weights based on impressions, or detectors applied to the same Beta–Binomial smoothed series used for the evaluation target–are natural extensions and should be evaluated before making strong robustness claims.
C.6 Computational Complexity Analysis
Lemma 1 (Per-window signature computation complexity (order of growth)).